# Geometry of the moduli space of Hermitian-Einstein connections on manifolds with a dilaton

### v1

@@ 2.2.1 Metric and Hermitian form on the moduli space
So far, the geometry of the Hermitian-Einstein connection moduli spaces has been investigated in the Gauduchon gauge, i.e. in the gauge that the Lee form $\theta$ of the metric on $M^{2n}$ is divergence free $D^{i}\theta_{i}=0$. This is not a restriction on the geometry of $(M^{2n},g,\omega)$ as the conformal class of every Hermitian metric $g$ admits such a metric. However, in many applications, especially in the context of string theory and in manifolds with special holonomy, the relevant metric on $(M^{2n},g,\omega)$ is not in the Gauduchon gauge and so $D^{i}\theta_{i}\not=0$. We shall overcome this difficulty by introducing a scalar field $\Phi$ on $M^{2n}$, the dilaton. As a result, we shall investigate the geometry of the moduli space of Hermitian-Einstein connections in a way that incorporates the dilaton $\Phi$. Luckily, this can be done as an adaptation of the results that have already derived in the Gauduchon gauge. As we shall point out, there is a class of geometries that such a scalar field naturally arises and these are precisely those that appear in the context of string theory, manifolds with special holonomy and Hermitian geometric flow solitons.
To begin consider a Hermitian manifold $(M^{2n},g,\omega)$ with metric $g$ and form $\omega$, where $g$ may not be in the Gauduchon gauge. Suppose that $M^{2n}$ also admits a scalar function $\Phi$. The Hermitian-Einstein condition (2.7) is taken with respect to the complex structure $I$ and Hermitian form $\omega$ of $M^{2n}$. As a result, the tangent vectors $a$ of $\mathscr{A}^{*}_{\mathrm{HE}}$ satisfy the conditions (2.8) but with a Hermitian form $\omega$ whose associated metric $g$ is not Gauduchon. Moreover for $\mathscr{M}_{\mathrm{HE}}$ to admit a complex structure, the gauge fixing, or horizontality condition, should be chosen as in (2.10). Clearly, the conditions on the tangent vectors of $\mathscr{A}^{*}_{\mathrm{HE}}$ to be along $\mathscr{A}^{*}_{\mathrm{HE}}$ and horizontal are the same as those we have investigated so far. This potentially can allow as to decompose the tangent vectors $a$ of $\mathscr{A}^{*}_{\mathrm{HE}}$ into horizontal $a^{h}$ and vertical components $a^{v}$ as
$a=a^{h}+a^{v}\leavevmode\nobreak\ ,$ (2.15)
@@
One way to incorporate the dilaton in the analysis of the geometry on the moduli spaces is to modify the inner product, and consequently the Hermitian form, on $\mathscr{A}$ as
$(a_{1},a_{2})_{\mathscr{A}}$ $\equiv\int_{M^{2n}}d^{n}x\sqrt{g}\,\,e^{-2\Phi}\,\,g^{-1}\,\langle a_{1},a_{2}\rangle_{\mathfrak{g}}=\int_{M^{2n}}d^{n}x\sqrt{g}\,\,e^{-2\Phi}\,\,g^{ij}\,\langle a_{1i},a_{2j}\rangle_{\mathfrak{g}}\leavevmode\nobreak\ ,$ (2.16)
$\Omega_{\mathscr{A}}(a_{1},a_{2})$ $\equiv(a_{1},\mathcal{I}a_{2})_{\mathscr{A}}=\frac{1}{(n-1)!}\,\int_{M^{2n}}\,\,e^{-2\Phi}\,\,\omega^{n-1}\wedge\langle a_{1}\wedge a_{2}\rangle_{\mathfrak{g}}\leavevmode\nobreak\ .$ (2.17)
Assuming that the decomposition (2.15) is unique, the inner product and Hermitian form on the moduli space $\mathscr{M}^{*}_{\mathrm{HE}}$ can be defined as
$(\alpha_{1},\alpha_{2})_{\mathscr{M}^{*}_{\mathrm{HE}}}$ $\equiv(a^{h}_{1},a^{h}_{2})_{\mathscr{A}^{*}_{\mathrm{HE}}}\leavevmode\nobreak\ ,$ (2.18)
$\Omega_{\mathscr{M}^{*}_{\mathrm{HE}}}(\alpha_{1},\alpha_{2})$ $\equiv\Omega_{\mathscr{A}^{*}_{\mathrm{HE}}}(a^{h}_{1},a^{h}_{2})\leavevmode\nobreak\ ,$ (2.19)
@@
Let $(M^{2n},g,\omega)$ be a Hermitian manifold with dilaton field $\Phi$ and Lee form $\theta$. The decomposition of tangent vector $a$ on $\mathscr{A}^{*}_{\mathrm{HE}}$ in horizontal and vertical components as $a=a^{h}+a^{v}$ is unique provided that
$D^{i}\big{(}e^{-2\Phi}K^{\flat}_{i}\big{)}=\big{(}D^{i}(e^{-2\Phi}\theta_{i})-D^{2}e^{-2\Phi}\big{)}=0\leavevmode\nobreak\ ,$ (2.20)
@@
$D_{A}^{i}a_{i}+\theta^{i}a_{i}=(D^{i}_{A}D_{Ai}+\theta^{i}D_{Ai})\epsilon\leavevmode\nobreak\ ,$ (2.21)
@@
$\mathcal{O}=D^{i}_{A}D_{Ai}+\theta^{i}D_{Ai}\leavevmode\nobreak\ ,$ (2.22)
@@
$\qquad=-\int_{M}d^{2n}x\,\sqrt{g}\,\Big{(}e^{-2\Phi}\langle d_{A}^{i}\epsilon,(d_{A})_{i}\epsilon\rangle_{\mathfrak{g}}+D_{i}\big{(}e^{-2\Phi}(\partial^{i}\Phi+{1\over 2}\theta^{i})\big{)}\langle\epsilon,\epsilon\rangle_{\mathfrak{g}}\Big{)}=0\leavevmode\nobreak\ .$ (2.25)
Using the hypothesis of the theorem, the above identity implies that $d_{A}\epsilon=0$. As for irreducible connection $d_{A}\epsilon=0$ implies that $\epsilon=0$, the kernel of $\mathcal{O}$ is trivial.
@@
$=\int_{M}\sqrt{g}d^{2n}xe^{-2\Phi}\langle\big{(}D_{A}^{i}D_{Ai}-(\theta_{i}+4\partial_{i}\Phi)D_{A}^{i}\big{)}\eta,\epsilon\rangle_{\mathfrak{g}}=0\leavevmode\nobreak\ ,$ (2.27)
@@
The condition (2.20) is satisfied provided that the 1-form $K^{\flat}=\theta+2d\Phi$ is Killing and $K\Phi=0$. As we shall demonstrate this condition allows for the inclusion in this analysis of the geometries that occur in string theory, those on manifolds with special holonomy and those of solitons of Hermitian flows. Moreover, (2.20) can possibly be weaken further, e.g if $D^{i}\big{(}e^{-2\Phi}(\partial_{i}\Phi+{1\over 2}\theta_{i})\big{)}\geq 0$, the kernel of $\mathcal{O}$ is trivial. However in what follows, the stronger version of the condition stated in the theorem above will suffice.
@@ 2.2.2 Geometry of moduli spaces
Before we proceed with the proof of one of the main theorems, let $a_{2}^{h}$ be a horizontal vector field $D_{A}^{i}a_{2i}+\theta^{i}a_{2i}=0$. Taking the directional derivative [^13] The directional derivative of a function $f$ on $\mathscr{A}$ along the vector field $a$ is defined as $a\cdot f(A)\equiv\frac{d}{dt}f(A+ta)|_{t=0}$. The commutator of two vector fields $a_{1}$ and $a_{2}$ on $\mathscr{A}$ is then defined as $[\kern-2.0pt[a_{1},a_{2}]\kern-2.0pt]f\equiv a_{1}\cdot(a_{2}\cdot f)-a_{2}\cdot(a_{1}\cdot f)$. of this condition along $a^{h}_{1}$ and then the commutator of $a^{h}_{1}$ and $a^{h}_{2}$, one finds that
@@
where $[\kern-2.0pt[a_{1},a_{2}]\kern-2.0pt]^{v}$ is the vertical component of the commutator. As this is a vector field along the orbits of the gauge group, there is a $\Theta$ such that $[\kern-2.0pt[a_{1},a_{2}]\kern-2.0pt]^{v}=-d_{A}\Theta(a_{1}^{h},a_{2}^{h})$ with
$\mathcal{O}\Theta(a_{1}^{h},a_{2}^{h})=2g^{ij}[a_{1i},a_{2j}]_{\mathfrak{g}}\leavevmode\nobreak\ .$ (2.29)
$\Theta$ is interpreted as the curvature associated with the horizontality condition (2.10). As expected it is a function on $M^{2n}$ and a 2-form on $\mathscr{M}^{*}_{\mathrm{HE}}$. The proof of the theorem below is structured along the lines of that presented in [27] for Gauduchon manifolds.
@@
$d\Omega_{\mathscr{M}^{*}_{\mathrm{HE}}}$ $(\alpha_{1},\alpha_{2},\alpha_{3})=-\frac{1}{(n-1)!}\,\int_{M^{2n}}\,\,e^{-2\Phi}\,\,\omega^{n-1}\wedge\langle d_{A}\Theta(a_{1}^{h},a_{2}^{h})\wedge a^{h}_{3}\rangle_{\mathfrak{g}}+\mathrm{cyclic\leavevmode\nobreak\ in\leavevmode\nobreak\ }(a^{h}_{1},a^{h}_{2},a^{h}_{3})$ (2.30)
$=\frac{1}{(n-1)!}\,\int_{M^{2n}}d(e^{-2\Phi}\,\,\omega^{n-1})\wedge\langle\Theta(a_{1}^{h},a_{2}^{h}),a^{h}_{3}\rangle_{\mathfrak{g}}+\mathrm{cyclic\leavevmode\nobreak\ in\leavevmode\nobreak\ }(a^{h}_{1},a^{h}_{2},a^{h}_{3})$ (2.31)
@@
$=-\frac{1}{(n-1)!}\,\int_{M^{2n}}d^{c}\big{(}e^{-2\Phi}\,\,\omega^{n-1}\big{)}\wedge\Big{(}\langle\Theta(a_{1}^{h},a_{2}^{h}),a^{h}_{3}\rangle_{\mathfrak{g}}+\mathrm{cyclic\leavevmode\nobreak\ in\leavevmode\nobreak\ }(a^{h}_{1},a^{h}_{2},a^{h}_{3})\Big{)}$ (2.33)
$=-\,\int_{M^{2n}}d^{2n}x\,\sqrt{g}\,\,e^{-2\Phi}\,\,(\theta_{i}+2\partial_{i}\Phi)\,\,g^{ij}\,\Big{(}\langle\Theta(a_{1}^{h},a_{2}^{h}),a^{h}_{3j}\rangle_{\mathfrak{g}}+\mathrm{cyclic\leavevmode\nobreak\ in\leavevmode\nobreak\ }(a^{h}_{1},a^{h}_{2},a^{h}_{3})\Big{)}\leavevmode\nobreak\ .\leavevmode\nobreak\ .$ (2.34)
@@
$\qquad\qquad\qquad+\mathrm{cyclic\leavevmode\nobreak\ in\leavevmode\nobreak\ }(a^{h}_{1},a^{h}_{2},a^{h}_{3})\Big{)}$ (2.37)
@@
$\qquad\qquad\qquad+\mathrm{cyclic\leavevmode\nobreak\ in\leavevmode\nobreak\ }(a^{h}_{1},a^{h}_{2},a^{h}_{3})\Big{)}=0\leavevmode\nobreak\ ,$ (2.39)
@@
Conformally balanced is a class of Hermitian manifolds $(M^{2n},g,\omega)$ for which the Lee form is an exact 1-form, i.e. $\theta=-2d\Phi$. Clearly for those, $\mathscr{M}^{*}_{\mathrm{HE}}$ is a Kähler manifold – observe that the torsion ${\mathcal{H}}$ in (2.34) vanishes . Such Hermitian manifolds occur in supersymmetric string compactifications – the conformal balanced condition is a requirement imposed on the geometry of the internal space $M^{2n}$ as a consequence of the dilatino Killing spinor equation. Moreover, the gaugino Killing spinor equation implies that the connection of the gauge sector satisfies the Hermitian-Einstein condition. Therefore, one consequence of our results is that for fixed metric and complex structure on the internal space, the part of the moduli space of such compactifications due to the choice of a Hermitian-Einstein connection is a Kähler manifold, see also [9, 10, 11]. It should be also noted that the conformal balanced condition puts strong restrictions on Hermitian manifolds. For example, compact conformally balanced Hermitian manifolds $(M^{2n},g,\omega)$ with a strong KT structure, $dH=0$, whose holonomy of the $\widehat{\nabla}$ connection is included in $SU(n)$ are Calabi-Yau and so $H=0$ [47].

### v2

@@ 2.2.1 Metric and Hermitian form on the moduli space
So far, the geometry of the Hermitian-Einstein connection moduli spaces has been investigated in the Gauduchon gauge, i.e. in the gauge that the Lee form $\theta$ of the metric on $M^{2n}$ is divergence free $D^{i}\theta_{i}=0$. This is not a restriction on the geometry of $(M^{2n},g,\omega)$ as the conformal class of every Hermitian metric $g$ admits such a metric. However, in many applications, especially in the context of string theory and in manifolds with special holonomy, the relevant metric on $(M^{2n},g,\omega)$ is not in the Gauduchon gauge and so $D^{i}\theta_{i}\not=0$. We shall overcome this difficulty by introducing a scalar field $\Phi$ on $M^{2n}$, the dilaton. As a result, we shall investigate the geometry of the moduli space of Hermitian-Einstein connections in a way that incorporates the dilaton $\Phi$. Luckily, this can be done as an adaptation of the results that have already derived in the Gauduchon gauge. As we shall point out, there is a class of geometries that such a scalar field naturally arises and these are precisely those that appear in the context of string theory, manifolds with special holonomy and geometric flow gradient (generalised) solitons.
To begin consider a Hermitian manifold $(M^{2n},g,\omega)$ with metric $g$ and form $\omega$, where $g$ may not be in the Gauduchon gauge. Suppose that $M^{2n}$ also admits a scalar function $\Phi$. The Hermitian-Einstein condition (2.7) is taken with respect to the complex structure $I$ and Hermitian form $\omega$ of $M^{2n}$. As a result, the tangent vectors $a$ of $\mathscr{A}^{*}_{\mathrm{HE}}$ satisfy the conditions (2.8) but with a Hermitian form $\omega$ whose associated metric $g$ is not Gauduchon. Moreover for $\mathscr{M}_{\mathrm{HE}}$ to admit a complex structure, the gauge fixing, or horizontality condition, should be chosen as in (2.10), again with respect to the metric $g$ (and Hermitian form $\omega$) that do not satisfy the Gauduchon condition. Clearly, the conditions on the tangent vectors of $\mathscr{A}^{*}_{\mathrm{HE}}$ to be along $\mathscr{A}^{*}_{\mathrm{HE}}$ and horizontal are the same as those we have investigated so far. This potentially can allow us to decompose the tangent vectors $a$ of $\mathscr{A}^{*}_{\mathrm{HE}}$ into horizontal $a^{h}$ and vertical components $a^{v}$ as
$a=a^{h}+a^{v}\penalty 10000\ ,$ (2.15)
@@
One way to incorporate the dilaton in the analysis of the geometry on the moduli spaces on non-Gauduchon Hermitian manifolds $(M^{2n},g,\omega)$ is to modify the inner product, and consequently the Hermitian form, on $\mathscr{A}$ as
$(a_{1},a_{2})_{\mathscr{A}}$ $\equiv\int_{M^{2n}}d^{n}x\sqrt{g}\,\,e^{-2\Phi}\,\,g^{-1}\,\langle a_{1},a_{2}\rangle_{\mathfrak{g}}=\int_{M^{2n}}d^{n}x\sqrt{g}\,\,e^{-2\Phi}\,\,g^{ij}\,\langle a_{1i},a_{2j}\rangle_{\mathfrak{g}}\penalty 10000\ ,$ (2.16)
$\Omega_{\mathscr{A}}(a_{1},a_{2})$ $\equiv(a_{1},\mathcal{I}a_{2})_{\mathscr{A}}=\frac{1}{(n-1)!}\,\int_{M^{2n}}\,\,e^{-2\Phi}\,\,\omega^{n-1}\wedge\langle a_{1}\wedge a_{2}\rangle_{\mathfrak{g}}\penalty 10000\ ,$ (2.17)
see [27] for the notation in the last equation. Assuming that the decomposition (2.15) is unique, the inner product and Hermitian form on the moduli space $\mathscr{M}^{*}_{\mathrm{HE}}$ can be defined as
$(\alpha_{1},\alpha_{2})_{\mathscr{M}^{*}_{\mathrm{HE}}}$ $\equiv(a^{h}_{1},a^{h}_{2})_{\mathscr{A}^{*}_{\mathrm{HE}}}\penalty 10000\ ,$ (2.18)
$\Omega_{\mathscr{M}^{*}_{\mathrm{HE}}}(\alpha_{1},\alpha_{2})$ $\equiv\Omega_{\mathscr{A}^{*}_{\mathrm{HE}}}(a^{h}_{1},a^{h}_{2})\penalty 10000\ ,$ (2.19)
@@
Let $(M^{2n},g,\omega)$ be a (non-Gauduchon) Hermitian manifold with dilaton field $\Phi$ and Lee form $\theta$. The decomposition of tangent vector $a$ on $\mathscr{A}^{*}_{\mathrm{HE}}$ in horizontal and vertical components as $a=a^{h}+a^{v}$ is unique provided that
$D^{i}\big(e^{-2\Phi}K^{\flat}_{i}\big)=\big(D^{i}(e^{-2\Phi}\theta_{i})-D^{2}e^{-2\Phi}\big)=0\penalty 10000\ ,$ (2.20)
@@
$D_{A}^{i}a_{i}+\theta^{i}a_{i}=(D^{i}_{A}D_{Ai}+\theta^{i}D_{Ai})\epsilon\penalty 10000\ ,$ (2.21)
@@
$\mathcal{O}=D^{i}_{A}D_{Ai}+\theta^{i}D_{Ai}\penalty 10000\ ,$ (2.22)
@@
$\qquad=-\int_{M}d^{2n}x\,\sqrt{g}\,\Big(e^{-2\Phi}\langle d_{A}^{i}\epsilon,(d_{A})_{i}\epsilon\rangle_{\mathfrak{g}}+D_{i}\big(e^{-2\Phi}(\partial^{i}\Phi+{1\over 2}\theta^{i})\big)\langle\epsilon,\epsilon\rangle_{\mathfrak{g}}\Big)=0\penalty 10000\ .$ (2.25)
Using the hypothesis of the theorem, the above identity implies that $d_{A}\epsilon=0$. As for irreducible connections AA $d_{A}\epsilon=0$ implies that $\epsilon=0$, the kernel of $\mathcal{O}$ is trivial.
@@
$=\int_{M}\sqrt{g}d^{2n}x\,e^{-2\Phi}\langle\big(D_{A}^{i}D_{Ai}-(\theta_{i}+4\partial_{i}\Phi)D_{A}^{i}\big)\eta,\epsilon\rangle_{\mathfrak{g}}=0\penalty 10000\ ,$ (2.27)
@@
The condition (2.20) is satisfied provided that the 1-form $K^{\flat}=\theta+2d\Phi$ is Killing and $K\Phi=0$. As we shall demonstrate this condition allows for the inclusion in this analysis of the geometries that occur in string theory, those on manifolds with special holonomy and those of (generalised) solitons of Hermitian geometric flows. Moreover, (2.20) can possibly be weaken further, e.g if $D^{i}\big(e^{-2\Phi}(\partial_{i}\Phi+{1\over 2}\theta_{i})\big)\geq 0$, the kernel of $\mathcal{O}$ is trivial. However in what follows, the stronger version of the condition stated in the theorem above will suffice.
@@ 2.2.2 Geometry of moduli spaces
Before we proceed with the proof of one of the main theorems, let $a_{2}^{h}$ be a horizontal vector field $D_{A}^{i}a_{2i}+\theta^{i}a_{2i}=0$. Taking the directional derivative [^13] The directional derivative of a function $f$ on $\mathscr{A}$ along the vector field $a$ is defined as $a\cdot f(A)\equiv\frac{d}{dt}f(A+ta)|_{t=0}$. The commutator of two vector fields $a_{1}$ and $a_{2}$ on $\mathscr{A}$ is then defined as $\kern-2.0pt[a_{1},a_{2}]\kern-2.0pt]f\equiv a_{1}\cdot(a_{2}\cdot f)-a_{2}\cdot(a_{1}\cdot f)$. of this condition along $a^{h}_{1}$ and then the commutator of $a^{h}_{1}$ and $a^{h}_{2}$, one finds that
@@
where $\kern-2.0pt[a_{1},a_{2}]\kern-2.0pt]^{v}$ is the vertical component of the commutator. As this is a vector field along the orbits of the gauge group, there is a $\Theta$ such that $\kern-2.0pt[a^{h}_{1},a^{h}_{2}]\kern-2.0pt]^{v}=-d_{A}\Theta(a_{1}^{h},a_{2}^{h})$ with
$\mathcal{O}\Theta(a_{1}^{h},a_{2}^{h})=2g^{ij}[a^{h}_{1i},a^{h}_{2j}]_{\mathfrak{g}}\penalty 10000\ .$ (2.29)
$\Theta$ is interpreted as the curvature associated with the horizontality condition (2.10). As expected it is a Lie algebra valued 2-form on $\mathscr{M}^{*}_{\mathrm{HE}}$, where the Lie algebra is that of $\mathscr{G}^{*}$. The proof of the theorem below is structured along the lines of that presented in [27] for Gauduchon manifolds.
@@
$d\Omega_{\mathscr{M}^{*}_{\mathrm{HE}}}$ $(\alpha_{1},\alpha_{2},\alpha_{3})=-\frac{1}{(n-1)!}\,\int_{M^{2n}}\,\,e^{-2\Phi}\,\,\omega^{n-1}\wedge\langle d_{A}\Theta(a_{1}^{h},a_{2}^{h})\wedge a^{h}_{3}\rangle_{\mathfrak{g}}+\mathrm{cyclic\penalty 10000\ in\penalty 10000\ }(a^{h}_{1},a^{h}_{2},a^{h}_{3})$ (2.30)
$=\frac{1}{(n-1)!}\,\int_{M^{2n}}d(e^{-2\Phi}\,\,\omega^{n-1})\wedge\langle\Theta(a_{1}^{h},a_{2}^{h}),a^{h}_{3}\rangle_{\mathfrak{g}}+\mathrm{cyclic\penalty 10000\ in\penalty 10000\ }(a^{h}_{1},a^{h}_{2},a^{h}_{3})$ (2.31)
@@
$=-\frac{1}{(n-1)!}\,\int_{M^{2n}}d^{c}\big(e^{-2\Phi}\,\,\omega^{n-1}\big)\wedge\Big(\langle\Theta(a_{1}^{h},a_{2}^{h}),a^{h}_{3}\rangle_{\mathfrak{g}}+\mathrm{cyclic\penalty 10000\ in\penalty 10000\ }(a^{h}_{1},a^{h}_{2},a^{h}_{3})\Big)$ (2.33)
$=-\,\int_{M^{2n}}d^{2n}x\,\sqrt{g}\,\,e^{-2\Phi}\,\,(\theta_{i}+2\partial_{i}\Phi)\,\,g^{ij}\,\Big(\langle\Theta(a_{1}^{h},a_{2}^{h}),a^{h}_{3j}\rangle_{\mathfrak{g}}+\mathrm{cyclic\penalty 10000\ in\penalty 10000\ }(a^{h}_{1},a^{h}_{2},a^{h}_{3})\Big)\penalty 10000\ .\penalty 10000\ .$ (2.34)
@@
$\qquad\qquad\qquad+\mathrm{cyclic\penalty 10000\ in\penalty 10000\ }(a^{h}_{1},a^{h}_{2},a^{h}_{3})\Big)$ (2.37)
@@
$\qquad\qquad\qquad+\mathrm{cyclic\penalty 10000\ in\penalty 10000\ }(a^{h}_{1},a^{h}_{2},a^{h}_{3})\Big)=0\penalty 10000\ ,$ (2.39)
@@
Conformally balanced is a class of Hermitian manifolds $(M^{2n},g,\omega)$ for which the Lee form is an exact 1-form, i.e. $\theta=-2d\Phi$. Clearly for those, $\mathscr{M}^{*}_{\mathrm{HE}}$ is a Kähler manifold – observe that the torsion $\mathcal{H}$} in (2.34) vanishes . Such Hermitian manifolds occur in supersymmetric string compactifications – the conformal balanced condition is a requirement imposed on the geometry of the internal space $M^{2n}$ as a consequence of the dilatino Killing spinor equation. Moreover, the gaugino Killing spinor equation implies that the connection of the gauge sector satisfies the Hermitian-Einstein condition. Therefore, one consequence of our results is that for fixed metric and complex structure on the internal space, the part of the moduli space of such compactifications due to the choice of a Hermitian-Einstein connection is a Kähler manifold, see also [9, 10, 11]. It should be also noted that the conformal balanced condition puts strong restrictions on Hermitian manifolds. For example, compact conformally balanced Hermitian manifolds $(M^{2n},g,\omega)$ with a strong KT structure, $dH=0$, whose holonomy of the $\widehat{\nabla}$ connection is included in $SU(n)$, are Calabi-Yau and so $H=0$ [48].